# Tensor Product Attention Is All You Need

### v3

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, significantly shrinking KV cache size at inference time. By factorizing these representations into contextual low-rank components (contextual factorization) and seamlessly integrating with RoPE, TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation of language modeling tasks, we demonstrate that T6 exceeds the performance of standard Transformer baselines including MHA, MQA, GQA, and MLA across various metrics, including perplexity and a range of renowned evaluation benchmarks. Notably, TPA’s memory efficiency enables the processing of significantly longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. The code is available at https://github.com/tensorgi/T6.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completions. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (Zhang et al., 2023c; Liu et al., 2024c). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
Figure 1: Tensor Product Attention (TPA) in the Tensor ProducT ATTenTion Transformer (T6). Different from multi-head attention, in each layer, firstly the hidden state goes through different linear layers to get the latent factor matrices $\mathbf{A}_{(\cdot)}$’s and $\mathbf{B}_{(\cdot)}$’s for query, key, and value. We additionally apply RoPE to $\mathbf{B}_{Q}$ and $\mathbf{B}_{K}$ for query and key. Then the multi-head query, key, and value vectors are attained by the tensor product of $\mathbf{A}_{(\cdot)}$ and $\mathbf{B}_{(\cdot)}$. Finally, the output of TPA is produced by scaled dot-product attention followed by linear projection of concatenated results of multiple heads.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (He & Zhai, 2024), at the expense of increased I/O latency. Attention variants like multi-query attention (MQA) (Shazeer, 2019) and grouped-query attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates runtime. The recently introduced Multi-head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but needs additional position-encoded parameters per head due to incompatibility with Rotary Position Embedding (RoPE) efficiently (Su et al., 2024).
@@
In order to overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), as illustrated in Figure 1, a novel architecture that uses higher-order tensors to factorize queries (Q), keys (K), and values (V) during attention computation. By dynamically factorizing activations rather than static weights (e.g., LoRA), TPA constructs low-rank, contextual representations that substantially reduce KV cache memory usage with improved representational capacity. In practice, TPA can reduce the memory overhead by an order of magnitude compared to standard multi-head attention (MHA) with lower pretraining validation loss (perplexity) and improved downstream performance.
@@
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023) and Gemma (Team et al., 2024).
Our primary contributions are summarized as follows:
@@
  • We propose Tensor Product Attention (TPA), A mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor-decompositions to achieve 10×10$\times$10 × or more reduction in inference-time KV cache size relative to standard attention mechanism (Vaswani et al., 2017) with improved performance compared to previous methods such as MHA, MQA, GQA, and MLA. In addition, we unify existing attention mechanisms by revealing that MHA, MQA, and GQA all arise naturally as non-contextual variants of TPA.
@@
  • We introduce the Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling. On language modeling experiments, T6 consistently improves validation perplexity and downstream evaluation performance with reduced KV cache size.
@@
  • We show TPA integrates seamlessly with RoPE (Su et al., 2024), facilitating easy adoption in popular foundation model architectures such as LLaMA and Gemma.
@@ 2 Background
In this section, we review two classical forms of attention: Scaled Dot-Product Attention, and Multi-Head Attention (MHA) (Vaswani et al., 2017). More types of attention are detailed in the Appendix E, including Multi-Query Attention (MQA) (Shazeer, 2019), and Grouped Query Attention (GQA) (Ainslie et al., 2023), as well as a recent method called Multi-head Latent Attention (MLA) used in DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b). We also introduce Rotary Position Embedding (RoPE, Su et al. (2024)), which is commonly used in recent works of large language models.
@@
Notations. We use bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Q}$) for matrices, bold lowercase (e.g., $\mathbf{a}$, $\mathbf{b}$) for vectors, and italic uppercase (e.g., $\bm{W}^{Q}_{i}$) for learnable parameter matrices. We denote by $[n]$ the set $\{1,\ldots,n\}$ for some positive integer $n$. We use $\top$ to denote the transpose of a vector or a matrix. Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens. With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token.
Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively. In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head. [^5] Often, one sets $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$. Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$. For methods like MQA and GQA, some of these are shared or partially shared across heads, but their shapes remain consistent.
We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is:
$\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~{}C_{ij}=a_{i}b_{j},$
where $a_{i}$ and $b_{j}$ are the $i$-th and $j$-th elements of $\mathbf{a}$ and $\mathbf{b}$ respectively, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$. We also define the vectorization of a matrix $\mathbf{C}\in\mathbb{R}^{m\times n}$ by:
$\text{vec}(\mathbf{C})=\mathbf{d}\in\mathbb{R}^{mn},\text{with}~{}d_{i\cdot n+j}=C_{ij},$
where $d_{i\cdot n+j}$ is the $(i\cdot n+j)$-th element of $\mathbf{d}$.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (Vaswani et al., 2017) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
@@
where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$. The division by $\sqrt{d_{k}}$ stabilizes training by controlling the scale of the inner products.
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
$\mathbf{Q}_{t,i}$ $=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
$\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
$\mathbf{V}_{t,i}$ $=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
@@
where $\bm{W}_{i}^{Q},\bm{W}_{i}^{K},\bm{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ are learnable projection matrices for the $i$-th head, $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$. After computing each head’s attention, the outputs are concatenated and mapped back to the original dimension via another matrix $\bm{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{model}}}$:
$\operatorname{MHA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\,\bm{W}^{O}.$
MHA can capture a rich set of dependencies while each head focuses on different subspaces.
2.3 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; Su et al., 2024) to encode positional information in the query/key vectors. Specifically, let $\mathrm{RoPE}_{t}$ denote the rotation operator $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ corresponding to the $t$-th position. $\mathbf{T}_{t}$ is a block-diagonal matrix, which consists of block-diagonal matrix$\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$, $j\in\{1,\cdots,d_{h}/2\}$, where $\{\theta_{j}\}$ are pre-defined frequency parameters, e.g., $\theta_{j}=1/10000^{2j/d_{h}}$. Then we define
$\operatorname{RoPE}\left(\mathbf{Q}_{t}\right)\triangleq\mathbf{Q}_{t}\mathbf{T}_{t},\quad\text{where }\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}.$
A fundamental property is that
$\mathbf{T}_{t}\,\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (2.1)
which ensures that relative positions $(t-s)$ are preserved, thereby providing a form of translation invariance in the rotary position embedding.

### v4

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time. By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks. Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. The code is available at https://github.com/tensorgi/T6.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language, ; chowdhery2022palm, ; touvron2023llama, ; bubeck2023sparks, ). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o, ; liu2024kivi, ). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
Figure 1: Tensor Product Attention (TPA) within the Tensor ProducT ATTenTion Transformer (T6). In each TPA layer, the input hidden state $\mathbf{x}_{t}$ is processed by linear layers to produce latent factor matrices for query (e.g., $\mathbf{A}_{Q}(\mathbf{x}_{t}),\mathbf{B}_{Q}(\mathbf{x}_{t})$), key (e.g., $\mathbf{A}_{K}(\mathbf{x}_{t}),\mathbf{B}_{K}(\mathbf{x}_{t})$), and value (e.g., $\mathbf{A}_{V}(\mathbf{x}_{t}),\mathbf{B}_{V}(\mathbf{x}_{t})$). Rotary Position Embedding (RoPE) is applied to the $\mathbf{B}_{Q}(\mathbf{x}_{t})$ and $\mathbf{B}_{K}(\mathbf{x}_{t})$ factors. The query, key, and value tensors for each attention head are then formed by the tensor product of these factor matrices (e.g., $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$). Finally, the TPA output is computed using scaled dot-product attention, followed by a linear projection of the concatenated results from all heads.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating, ) or token eviction strategies (zhang2023h2o, ; xiao2023efficient, ; ribar2023sparq, ), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (he2024fastdecode, ), at the expense of increased I/O latency. Attention variants like Multi-Query Attention (MQA) (shazeer2019fast, ) and Grouped-Query Attention (GQA) (ainslie2023gqa, ) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora, ) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime. The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek, ) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer, ) integration, necessitating additional position-encoded parameters per head.
@@
To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1. TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V). By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations. This approach substantially reduces KV cache memory usage while offering improved representational capacity. In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance.
@@
A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer, ), enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama, ) and Gemma (team2024gemma, ).
Our main contributions are summarized as follows:
  1. 1.
@@
    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions. This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention, ), MHA, MQA, GQA, and MLA, while also improving performance. In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
  2. 2.
@@
    We introduce the Tensor ProducT ATTenTion Transformer (T6), a new TPA-based model architecture for sequence modeling. In language modeling experiments, T6 consistently improves or matches validation perplexity and downstream evaluation performance, all while maintaining a reduced KV cache size.
  3. 3.
@@
    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer, ), facilitating its easy adoption in popular foundation model architectures like LLaMA and Gemma.
  4. 4.
    We develop FlashTPA Decoding, an efficient autoregressive inference algorithm for TPA. Our empirical results show that FlashTPA Decoding can be faster than optimized MHA, MQA, GQA, and MLA decoding methods, particularly for long sequences.
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention, ) and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast, ), Grouped Query Attention (GQA) (ainslie2023gqa, ), Multi-head Latent Attention (MLA) (liu2024deepseek, ; liu2024deepseekv3, ), and Rotary Position Embedding (RoPE) (su2024roformer, ) are further discussed in Appendix H.
@@
Notations. We use bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Q}$) for matrices, bold lowercase (e.g., $\mathbf{a}$, $\mathbf{b}$) for vectors, and italic uppercase (e.g., $\bm{W}^{Q}_{i}$) for learnable parameter matrices. We denote by $[n]$ the set $\{1,\ldots,n\}$ for some positive integer $n$. We use $\top$ to denote the transpose of a vector or a matrix. Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens. With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token. Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively. In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head. [^5] Often, $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$. Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$. For methods like MQA and GQA, some of these projection matrices are shared or partially shared across heads, but their shapes remain consistent.
We define the tensor product of two vectors as follows: for vectors $\mathbf{a}\in\mathbb{R}^{m},\mathbf{b}\in\mathbb{R}^{n}$, the tensor product of $\mathbf{a}$ and $\mathbf{b}$ is: $\mathbf{a}\otimes\mathbf{b}=\mathbf{C}\in\mathbb{R}^{m\times n},\text{with}~{}C_{ij}=a_{i}b_{j}$, where $a_{i}$ is the $i$-th element of $\mathbf{a}$, $b_{j}$ is the $j$-th element of $\mathbf{b}$, and $C_{ij}$ is the $(i,j)$-th entry of $\mathbf{C}$. The vectorization of a matrix $\mathbf{C}\in\mathbb{R}^{m\times n}$, denoted $\text{vec}(\mathbf{C})\in\mathbb{R}^{mn}$, stacks the columns of $\mathbf{C}$ into a single column vector. For example, if $\mathbf{C}=[\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{n}]$ where $\mathbf{c}_{j}$ are columns, then $\text{vec}(\mathbf{C})=[\mathbf{c}_{1}^{\top},\mathbf{c}_{2}^{\top},\ldots,\mathbf{c}_{n}^{\top}]^{\top}$.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (vaswani2017attention, ) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
@@
where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$.
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) (vaswani2017attention, ) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
$\mathbf{Q}_{t,i}=(\bm{W}_{i}^{Q})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},~{}\mathbf{K}_{t,i}$ $=(\bm{W}_{i}^{K})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},~{}\mathbf{V}_{t,i}=(\bm{W}_{i}^{V})^{\top}\,\mathbf{x}_{t}\in\mathbb{R}^{d_{h}},$
@@
where $\bm{W}_{i}^{Q},\bm{W}_{i}^{K},\bm{W}_{i}^{V}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ are learnable projection matrices for the $i$-th head, and $\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ are the query, key, and value matrices for the $i$-th head over $T$ tokens. After computing each head’s attention output, the results are concatenated and mapped back to the model’s original dimension via another learnable linear projection matrix $\bm{W}^{O}\in\mathbb{R}^{hd_{h}\times d_{\text{model}}}$:
$\operatorname{MHA}(\mathbf{X})=\operatorname{Concat}\bigl{(}\textbf{head}_{1},\ldots,\textbf{head}_{h}\bigr{)}\,\bm{W}^{O}.$
MHA enables the model to capture a rich set of dependencies by allowing each head to focus on different aspects of the input sequence. We also discuss how MHA relates to TPA in Appendix A. Further discussion on Transformers and attention mechanisms, KV cache optimization techniques, and low-rank factorization methods is available in Appendix G.